# Efficient Adaptive Experimentation with Non-Compliance

### v1

@@ 1 Introduction
Adaptive experimentation enables efficient estimation of treatment effects in sequential settings by adjusting assignment strategies based on accumulating data. Compared to traditional randomized controlled trials (RCTs), adaptive designs can reduce estimation variance, accelerate discovery, and limit exposure to ineffective or costly interventions. These methods are now widely used across domains—from medicine to technology—and have been formally endorsed by the U.S. Food and Drug Administration [20], driving both practical adoption and theoretical advances [22, 25, 13].
In many important settings, however, direct treatment assignment is not feasible or ethical. In medical and behavioral studies, researchers may only be able to encourage behaviors (e.g., physical activity) rather than enforce them. Online platforms may launch digital campaigns to encourage user actions–such as trying new features or subscribing to services–but cannot compel engagement. In such cases, noncompliance arises: individuals self-select into treatment based on unobserved factors that also affect outcomes, rendering standard estimators biased due to confounding.
To address this challenge, researchers often use instrumental variables (IVs)— randomized assignments that affect treatment uptake but not outcomes directly. However, while IV-based methods are well-established in static settings, and adaptive designs have been developed when treatment is assignable [25, 26, 13], adaptive experimentation with IVs remains largely unexplored. This creates a critical need for methods that can adaptively assign instruments and estimate treatment effects under noncompliance and unobserved confounding.
@@
This paper addresses this gap. We consider the problem of estimating the average treatment effect (ATE) in a sequential experiment where the experimenter can assign only a binary instrument, while the binary treatment is determined endogenously. We build on the semiparametric framework of Wang and Tchetgen Tchetgen [44], which identifies the ATE under an unconfounded compliance assumption and provides a multiply robust, efficient influence-function–based estimator. Unlike prior IV methods—which often target the local average treatment effect (LATE), assume structural equation models (SEMs), or lack robustness and efficiency—this framework enables ATE identification and efficient estimation without strong modeling assumptions. However, existing results apply only to static designs with fixed policies. We extend this framework to the adaptive setting. Specifically:
@@
  • We derive the semiparametric efficiency bound and characterize the optimal adaptive policy that minimizes this bound through a covariate-dependent instrument assignment.
@@
  • We introduce AMRIV, an Adaptive, Multiply Robust estimator for IV settings, which applies a sequential, plug-in version of the efficient influence function evaluated under the learned policy.
@@
  • We provide strong theoretical guarantees, including asymptotic normality, convergence rates, multiply robust consistency, and time-uniform asymptotic confidence sequences for valid inference at arbitrary stopping times.
@@
  • We demonstrate practical effectiveness in both synthetic and semi-synthetic settings, showing improved efficiency and robustness over non-adaptive baselines and alternative methods.
@@ 2 Related Work
IV-Based Estimation of ATE. ATE identification in IV settings has traditionally relied on structural equation models (SEMs) that impose parametric assumptions on the outcome and treatment assignment mechanisms. More recent work has proposed flexible alternatives—such as DeepIV [23], kernel IV [40], and orthogonal moment methods [42, 6]—that enable conditional effect estimation in high-dimensional or nonlinear settings. However, these approaches do not directly target robustness or semiparametric efficiency for the ATE. We instead build on the framework of Wang and Tchetgen Tchetgen [44], which establishes point identification of the ATE via an unconfounded compliance assumption without requiring SEMs. Their influence-function–based estimator achieves semiparametric efficiency and is multiply robust, remaining consistent under partial nuisance misspecification. We extend this framework to the adaptive setting and use it as the foundation for our estimator.
Adaptive ATE Estimation. Recent work has established a rich theory for adaptive estimation of the ATE when treatments can be directly assigned. Kato et al. [25] introduced the A2IPW estimator for variance-optimal adaptive assignment; Kato et al. [26] extended this to settings with estimated policies and nuisance functions, and Cook et al. [13] further incorporated policy truncation and provided the first asymptotically valid confidence sequences. We integrate these advances—adaptive nuisance estimation, sample splitting, policy truncation, and time-uniform inference—into our estimator to enable efficient, robust learning in experiments with noncompliance.
Adaptive Experimentation with IVs. A small but growing literature explores adaptive design in settings where only an instrument, rather than the treatment, can be assigned. Closest to our work is Chandak et al. [9], who propose a practical influence-function–based procedure to reduce prediction error in nonparametric IV regression. However, they focus on prediction accuracy and do not address semiparametric efficiency or robustness. Other approaches focus on partial identification [3], adaptive data acquisition [21], or regret minimization in bandit-style settings with endogeneity [50, 16]. In contrast, our goal is to enable efficient and robust adaptive ATE estimation under noncompliance, and we are the first to construct an efficient, multiply robust estimator in this setting.
@@ 3 Background and Setup
We consider the problem of estimating the average treatment effect (ATE) of a binary treatment $A\in\{0,1\}$ on a real-valued outcome $Y\in\mathbb{R}$, in the presence of unobserved confounding, within an adaptive experimental setting. We adopt the potential outcomes framework, where each unit is associated with two potential outcomes, $Y(0)$ and $Y(1)$, corresponding to the outcomes under control and treatment, respectively. However, only the realized outcome $Y=Y(A)$ is ever observed.
@@
Because direct treatment assignment may be infeasible, we rely on a binary instrumental variable $Z\in\{0,1\}$ that influences treatment uptake. The instrument can be interpreted as a recommendation or encouragement—something the experimenter can assign, unlike the treatment itself. We use $Y(a,z)$ to denote the potential outcome under treatment $a\in\{0,1\}$ and instrument $z\in\{0,1\}$, and define $A(z)$ as the potential treatment that would be taken under instrument assignment $z\in\{0,1\}$.
@@
This allows the instrument-assignment policy to evolve over time based on accumulated data. Following the instrument assignment $Z_{t}$, the treatment $A_{t}=A(Z_{t})$ is realized, and the outcome $Y_{t}=Y(A_{t},Z_{t})$ is observed. The full observation at time $t$ is thus $(X_{t},Z_{t},A_{t},Y_{t})$. After $T$ rounds, the experimenter estimates the ATE from accumulated data $\mathcal{H}_{T}=\{(X_{i},Z_{i},A_{i},Y_{i})\}_{i=1}^{T}$.
@@
Notation: We write $\pi_{t}(X_{t}\mid\mathcal{H}_{t}):=\pi_{t}(1\mid X_{t},\mathcal{H}_{t})$ for the probability of assigning $Z_{t}=1$ given covariates and history. The $L_{2}$ norm of a function $f$ is $\|f\|_{2}:=\mathbb{E}_{P}[f(X)^{2}]^{1/2}$, and $\widehat{f}_{t}$ denotes an estimate of $f$ based on $t$ samples. We use $\widehat{\mathbb{E}}$ to denote empirical expectations computed from data.
@@ 4 Efficiency Bounds and Optimal Instrument Assignment
Proofs of Theorem 1 and Corollary 2 appear in Appendix C. Unlike standard adaptive ATE estimation, the optimal policy $\pi^{*}(X)$ prioritizes arms with higher residual variance, balancing both outcome and compliance noise to minimize estimator variance.
Connection to Neyman Allocation. In the special case of perfect compliance—i.e., when $A=Z$—the treatment is fully determined by the (conditionally) randomized instrument and our setting becomes the adaptive ATE estimation scenario. In this setting, $\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X)=\operatorname*{{\rm Var}}(Y-A\delta(X)\mid A=z,X)=\operatorname*{{\rm Var}}(Y\mid A=z,X)$ and thus the optimal allocation reduces to $\frac{\sqrt{\operatorname*{{\rm Var}}(Y\mid A=1,X)}}{\sqrt{\operatorname*{{\rm Var}}(Y\mid A=0,X)}+\sqrt{\operatorname*{{\rm Var}}(Y\mid A=1,X)}}$ which exactly matches the classical Neyman allocation for minimizing the variance of a difference-in-means estimator [33, 25]. This highlights that our policy generalizes Neyman allocation to settings with noncompliance and endogenous treatment, adjusting for both outcome and compliance-driven noise.
@@ 5.1 Adaptive Instrument Assignment
Residual–variance estimation. We estimate $\widehat{\sigma}^{2}_{t-1}(z,X)$ using the decomposition
@@
Unbiased two-stage estimation via cross-fitting. To mitigate finite-sample bias in estimating $\widehat{\sigma}^{2}_{t-1}(z,X)$, we apply the sequential cross-fitting scheme of Waudby-Smith et al. [45]. Thus, we split $\mathcal{H}_{t-1}$ into two temporal folds $\mathcal{H}_{t-1}^{(j)}=\{(X_{i},Z_{i},A_{i},Y_{i}):i\in[t-1],i\mod 2=j\},\;j\in\{0,1\}$, fit $\widehat{\delta}_{t-1}$ on one and compute residuals $\widehat{R}_{t-1}^{2}$ on the other, and vice-versa. The combined residuals are used to regress $\widehat{R}_{t-1}^{2}$ on $(Z,X)$ to estimate $\widehat{s}_{t-1}(z,X)$. Since $\widehat{\mu}^{Y}_{t-1}$ and $\widehat{\mu}^{A}_{t-1}$ do not depend on other nuisances, they are learned on the full history $\mathcal{H}_{t-1}$.
@@
where $k_{t}\in[2,\infty)$ is a truncation parameter satisfying $k_{t}\to\infty$ as $t\to\infty$. Truncation ensures that the instrument assignment probabilities remain bounded away from 00 and $1$, thereby improving finite-sample stability and leading to better theoretical guarantees.
@@ 5.2 AMRIV: Adaptive Multiply Robust Estimation of the ATE
Nuisance Estimation. The nuisance functions ${\widehat{\mu}^{Y}_{t-1}(0,X),\widehat{\mu}^{A}_{t-1}(0,X),\widehat{\delta}_{t-1}^{A}(X),\widehat{\delta}_{t-1}(X)}$ can be estimated using any flexible nonparametric regression method applied to the historical data $\mathcal{H}_{t-1}$. To reduce computational overhead, we can reuse the estimates of $\widehat{\mu}^{Y}_{t-1}(0,X)$ and $\widehat{\mu}^{A}_{t-1}(0,X)$ previously obtained for instrument assignment in Section 5.1. Similarly, the estimate of $\widehat{\delta}_{t-1}(X)$ can be formed by averaging the cross-fitted estimates $\widehat{\delta}_{t-1}^{(0)}$ and $\widehat{\delta}_{t-1}^{(1)}$, trained on the two data folds $\mathcal{H}_{t-1}^{(0)}$ and $\mathcal{H}_{t-1}^{(1)}$, respectively. The only remaining component is $\widehat{\delta}_{t-1}^{A}(X)$, which must be estimated separately if it was not already computed as part of the $\widehat{\delta}{t-1}(X)$ estimation pipeline.
@@
$\quad\cdot\left[Y_{t}-A_{t}\widehat{\delta}_{t-1}(X_{t})-\widehat{\mu}^{Y}_{t-1}(0,X_{t})+\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(X_{t})\right]+\widehat{\delta}_{t-1}(X_{t})$ (9)

### v2

# Efficient Adaptive Experimentation with Noncompliance

@@ 1 Introduction
Adaptive experimentation enables efficient estimation of treatment effects in sequential settings by adjusting assignment strategies based on accumulating data. Compared to traditional randomized controlled trials (RCTs), adaptive designs can reduce estimation variance, thus accelerating discovery and limiting exposure to ineffective or harmful interventions. These methods are now widely used across domains–from medicine to online platforms–and have been formally endorsed by the U.S. Food and Drug Administration [21], driving both practical adoption and theoretical advances.
In many such settings, however, direct treatment assignment is not feasible or ethical. The treatment must instead be encouraged through a randomized recommendation or design choice—often referred to as an instrumental variable (IV)—leaving the final decision to the participant. This gives rise to noncompliance, where the assigned encouragement and the actual treatment may differ, and where self-selection based on unobserved factors introduces confounding that biases standard estimators. For instance, in a real TripAdvisor experiment, users were exposed to different premium sign-up interfaces that encouraged membership enrollment [52]. The actual treatment—whether a user subscribed—could not be enforced, but the interface (the instrument) could be randomized or adaptively adjusted. Similarly, in clinical trials, a physician may recommend a new medication but cannot compel adherence: the recommendation can be assigned, yet treatment uptake remains endogenous. Related challenges arise in recommender systems and public health interventions, where engagement or behavioral uptake is voluntary.
In such applications, reducing estimator variance is not merely a statistical preference; it determines how quickly and confidently a study can reach conclusions. Because high variance delays both the detection of harmful effects and the confirmation of beneficial ones, adaptive designs that learn to allocate instruments in variance-minimizing ways can mitigate these risks, yielding tighter confidence sequences and enabling earlier, statistically valid stopping decisions. Despite extensive work on adaptive designs for settings where treatment can be directly enforced [23, 27, 13], adaptive experimentation under noncompliance—where treatment is voluntary but encouragement can be adaptively controlled, remains largely unexplored, even though it describes many real-world scenarios.
@@
This paper addresses this gap. We study the problem of estimating the average treatment effect (ATE) in a sequential experiment where the experimenter can assign only a binary instrument, while the treatment itself is determined endogenously. Building on the semiparametric framework of Wang and Tchetgen Tchetgen [54], which identifies the ATE under an unconfounded compliance assumption and provides a multiply robust, efficient influence-function–based estimator, we extend this framework to the adaptive setting. Specifically:
@@
  • We derive the semiparametric efficiency bound and characterize the variance-optimal adaptive policy that minimizes this bound through a covariate-dependent instrument assignment.
@@
  • We introduce AMRIV, an Adaptive, Multiply Robust estimator for IV settings, which applies a sequential, plug-in version of the efficient influence function evaluated under the adaptive policy.
@@
  • We establish strong theoretical guarantees, including asymptotic normality, explicit convergence rates, multiply robust consistency, and time-uniform asymptotic confidence sequences for valid inference at arbitrary stopping times.
@@
  • We demonstrate practical effectiveness in both synthetic and semi-synthetic studies, showing improved efficiency and robustness over non-adaptive baselines and alternatives.
@@ 2 Related Work
IV-Based ATE Estimation. ATE identification in IV settings has traditionally relied on structural equation models (SEMs) that impose parametric assumptions on the outcome and treatment assignment mechanisms. More recent work has proposed flexible alternatives—such as DeepIV [24], kernel IV [50], and orthogonal moment methods [52, 6]—that enable conditional effect estimation in high-dimensional or nonlinear settings. However, these approaches do not directly target robustness or semiparametric efficiency for the ATE. We instead build on the framework of Wang and Tchetgen Tchetgen [54], which establishes point identification of the ATE via an unconfounded compliance assumption without requiring SEMs. Their influence-function–based estimator achieves semiparametric efficiency and is multiply robust, remaining consistent under partial nuisance misspecification. We extend this framework to the adaptive setting and use it as the foundation for our estimator.
Adaptive Experimentation for ATE Estimation. A large and growing literature studies adaptive designs where the treatment itself can be directly assigned, with the goal of minimizing estimator variance or its regret analogue. Early two-stage designs asymptotically achieve the semiparametric efficiency bound [23], and fully sequential approaches such as A2IPW attain variance-optimal Neyman allocation [27]. Subsequent extensions learn the allocation policy online [28] and add principled policy truncation with the first anytime-valid confidence sequences [13]. Parallel work from an online-learning perspective achieves sublinear or logarithmic "Neyman regret" via clipped or optimistic algorithms [14, 38, 39], matches finite-sample lower bounds under low-switching policies [35], and extends to covariate-adaptive and off-policy settings [29, 33]. Together, these methods form a mature toolkit—adaptive nuisance learning, cross-fitting, policy truncation, regret-style allocation, and time-uniform inference—but all assume the experimenter can randomize the treatment directly. Our work generalizes these efficiency guarantees to the harder and less explored regime where only an instrument can be assigned and treatment uptake is endogenous.
Adaptive Experimentation with Instruments. A small but growing literature explores adaptive design in settings where only an instrument, rather than the treatment, can be assigned. Closest to our work is Chandak et al. [9], who propose a practical influence-function–based procedure to reduce prediction error in nonparametric IV regression. However, they focus on prediction accuracy and do not address semiparametric efficiency or robustness to nuisances. Other approaches focus on partial identification [3], adaptive data acquisition [22], or regret minimization in bandit-style settings with endogeneity [61, 16]. In contrast, our goal is to enable efficient and robust adaptive ATE estimation under noncompliance, providing the first efficient and multiply robust estimator for this setting.
@@ 3 Background and Setup
Figure 1: Left: Adaptive experimentation with noncompliance. Blue elements denote learned or assigned quantities ($\pi_{t}$, $Z_{t}$), orange elements represent observed variables ($X_{t}$, $A_{t}$, $Y_{t}$), and the dashed red arrow indicates noncompliance ($A_{t}(Z_{t})\neq Z_{t}$). Right: The adaptive policy yields faster confidence-sequence contraction, enabling earlier stopping.
We consider the problem of estimating the average treatment effect (ATE) of a binary treatment $A\in\{0,1\}$ on a real-valued outcome $Y\in\mathbb{R}$, in the presence of unobserved confounding, within an adaptive experimental setting. We adopt the potential outcomes framework, where each unit is associated with two potential outcomes, $Y(0)$ and $Y(1)$, corresponding to the outcomes under control and treatment, respectively. However, only the realized outcome $Y=Y(A)$, corresponding to the treatment actually received, is observed.
@@
Because direct treatment assignment may be infeasible, we rely on a binary instrumental variable $Z\in\{0,1\}$ that influences treatment uptake. The instrument can be interpreted as a recommendation or encouragement—something the experimenter can assign, unlike the treatment itself. We denote by $Y(a,z)$ the potential outcome under treatment level $a$ and instrument value $z$, and by $A(z)$ the potential treatment that would be taken under instrument assignment $z$. Noncompliance arises when the potential treatment does not follow the instrument, i.e., $A(z)\neq z$, reflecting endogenous treatment selection. Equivalently, a unit complies if its treatment adheres to the assigned instrument ($A(z)=z$).
@@
This allows the instrument-assignment policy to evolve over time based on accumulated data. Following the instrument assignment $Z_{t}$, the treatment $A_{t}=A_{t}(Z_{t})$ is realized, and the outcome $Y_{t}=Y(A_{t},Z_{t})$ is observed. The full observation at time $t$ is thus $(X_{t},Z_{t},A_{t},Y_{t})$. After $T$ rounds, the experimenter estimates the ATE from accumulated data $\mathcal{H}_{T}=\{(X_{i},Z_{i},A_{i},Y_{i})\}_{i=1}^{T}$.
@@
Remark 1 (Interpretation under violations of Assumption 2).
If Assumption˜2 does not hold, the ATE is no longer point-identified. The estimand then shifts to the average conditional local average treatment effect (ACLATE), which averages treatment effects among instrument-responsive individuals (compliers) across covariate strata. The ACLATE remains identified and interpretable, capturing how causal effects vary among compliers when compliance is confounded and cannot be fully controlled.
@@
Notation: We write $\pi_{t}(X_{t}\mid\mathcal{H}_{t}):=\pi_{t}(1\mid X_{t},\mathcal{H}_{t})$ for the probability of assigning $Z_{t}=1$ given covariates and history. The $L_{2}$ norm of a function $f$ is $\|f\|_{2}:=\mathbb{E}_{P}[f(X)^{2}]^{1/2}$, and $\widehat{f}_{t}$ denotes an estimate of $f$ based on $t$ samples. We use $\widehat{\mathbb{E}}$ to denote empirical expectations computed from data. Notation used throughout the paper is summarized in Appendix B (Table 1).
@@ 4 Efficiency Bounds and Optimal Instrument Assignment
Proofs of Theorem 1 and Corollary 2 are included in Appendix E.
Drivers of Efficient Allocation. From Corollary 2, the optimal assignment policy $\pi^{*}(X)$ allocates more weight to the arm $z$ with higher residual variance $\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X)$, where
$\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X)$ $=\operatorname*{{\rm Var}}(Y\mid Z=z,X)+\delta(X)^{2}\operatorname*{{\rm Var}}(A\mid Z=z,X)$
$\quad-2\delta(X)\operatorname*{Cov}(Y,A\mid Z=z,X).$
This expression reveals that, unlike in standard adaptive ATE estimation, the residual variance depends jointly on both outcome noise ($\operatorname*{{\rm Var}}(Y\mid Z=z,X)$) and compliance noise ($\operatorname*{{\rm Var}}(A\mid Z=z,X)$). When compliance is more uncertain in one arm, the estimator becomes noisier in that region, leading the optimal policy to allocate more probability mass to that arm to compensate.
Connection to Neyman Allocation. In the special case of perfect compliance (when $A(Z)=Z$), the treatment is fully determined by the (conditionally) randomized instrument and our setting becomes the classical adaptive ATE estimation scenario. In this setting, $\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X)=\operatorname*{{\rm Var}}(Y-A\delta(X)\mid A=z,X)=\operatorname*{{\rm Var}}(Y\mid A=z,X)$ and thus the optimal allocation reduces to $\frac{\sqrt{\operatorname*{{\rm Var}}(Y\mid A=1,X)}}{\sqrt{\operatorname*{{\rm Var}}(Y\mid A=0,X)}+\sqrt{\operatorname*{{\rm Var}}(Y\mid A=1,X)}}$ which exactly matches the classical Neyman allocation for minimizing the variance of a difference-in-means estimator [40, 27]. This highlights that our policy generalizes Neyman allocation to settings with noncompliance and endogenous treatment, adjusting for both outcome and compliance-driven noise.
@@ 5.1 Adaptive Instrument Assignment
Residual–variance estimation. One option for estimating $\widehat{\sigma}^{2}_{t-1}(z,X)$ is using the decomposition
@@
Unbiased two-stage estimation via cross-fitting. To mitigate finite-sample bias in estimating $\widehat{\sigma}^{2}_{t-1}(z,X)$, we apply the sequential cross-fitting scheme of Waudby-Smith et al. [55]. Thus, we split $\mathcal{H}_{t-1}$ into two temporal folds $\mathcal{H}_{t-1}^{(j)}=\{(X_{i},Z_{i},A_{i},Y_{i}):i\in[t-1],i\mod 2=j\},\;j\in\{0,1\}$, fit $\widehat{\delta}_{t-1}$ on one and compute residuals $\widehat{R}_{t-1}^{2}$ on the other, and vice-versa. The combined residuals are used to regress $\widehat{R}_{t-1}^{2}$ on $(Z,X)$ to estimate $\widehat{s}_{t-1}(z,X)$. Since $\widehat{\mu}^{Y}_{t-1}$ and $\widehat{\mu}^{A}_{t-1}$ do not depend on other nuisances, they can be learned on the full history $\mathcal{H}_{t-1}$.
@@
Remark 2 (Choice of nuisance and variance estimators).
Sequential cross-fitting removes the need for restrictive conditions (e.g., Donsker)—standard nonparametric convergence rates suffice for the guarantees in our main theorems. In practice, any consistent learner (e.g., $k$-NN, random forests, or neural networks) can be used, depending on data structure and sample size (see Appendix C). The variance estimator in Eq. (6) ensures nonnegativity via a small floor $\varepsilon$, though fully nonnegative alternatives such as self-normalized kernel estimators may also be used. Appendix C further discusses these options and outlines online or streaming implementations that avoid full data storage.
@@
where $k_{t}\in[2,\infty)$ is a truncation parameter satisfying $k_{t}\to\infty$ as $t\to\infty$. Truncation ensures that the instrument assignment probabilities remain bounded away from 0 and $1$, thereby improving finite-sample stability and leading to better theoretical guarantees.
@@ 5.2 AMRIV: Adaptive Multiply Robust Estimation of the ATE
Nuisance Estimation. The nuisance functions $\widehat{\mu}^{Y}_{t-1}(0,X),\widehat{\mu}^{A}_{t-1}(0,X),\widehat{\delta}_{t-1}^{A}(X),\widehat{\delta}_{t-1}(X)$} can be estimated using any flexible nonparametric regression method applied to the historical data $\mathcal{H}_{t-1}$. To reduce computational overhead, we can reuse the estimates of $\widehat{\mu}^{Y}_{t-1}(0,X)$ and $\widehat{\mu}^{A}_{t-1}(0,X)$ previously obtained for instrument assignment in Section 5.1. Similarly, the estimate of $\widehat{\delta}_{t-1}(X)$ can be formed by averaging the cross-fitted estimates $\widehat{\delta}_{t-1}^{(0)}$ and $\widehat{\delta}_{t-1}^{(1)}$, trained on the two data folds $\mathcal{H}_{t-1}^{(0)}$ and $\mathcal{H}_{t-1}^{(1)}$, respectively. The only remaining component is $\widehat{\delta}_{t-1}^{A}(X)$, which must be estimated separately if it was not already computed as part of the $\widehat{\delta}_{t-1}(X)$ estimation pipeline.
@@
$\qquad\cdot\left[Y_{t}-A_{t}\widehat{\delta}_{t-1}(X_{t})-\widehat{\mu}^{Y}_{t-1}(0,X_{t})+\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(X_{t})\right]+\widehat{\delta}_{t-1}(X_{t})$ (9)